Every NASCAR race car has a wheel diameter of 15 inches width and an additional 3 inches for tires. That means the radius would be 0.27m. During a race these cars travel at 200 mph (89.4m/s). What is the time for one rotation of a NASCAR wheel during a race?


1
Expert's answer
2021-03-18T14:02:31-0400
"T=\frac{2\pi r}{v}\\T=\frac{2\pi(0.27)}{89.4}=0.019\ s"